Mason Brewing sells craft beer and runs a brewery academy offering brewing lessons in
Vancouver. On June 30th, 2018, Mason Brewing purchased a new Copper Kettle to brew
a special new micro-brew.
Mason brewing incurred the following costs:
1) The final invoice price for the copper kettle was $ 35,340
2) Due to it's unusual size, shipping costs of $ 830 were incurred
3) Mason Brewing purchased a one year insurance policy costing $ 2,800 to mitigate risk
4) Mason brewing hired specialists from the supplier to complete installation of the kettle for $ 630
5) To ensure quality from the very first batch, Mason brewing spent $ 1,200 testing the equipment
Mason brewing estimates they can use the new copper kettle for 10 years, at which time they could
probably sell it to a discount brewery for $ 3,000 . They also expect to use the kettle to brew
approximately 900,000 litres of beer.
Required:
1) Record the entries for the costs incurred on June 30th, 2018 to acquire the Copper Kettle. Assume
all costs were paid in cash.

2) Record depreciation at December 31st 2018 AND December 31st, 2019 assuming Mason Brewing uses
the following depreciation methods
A) Straight-line
B) Units of production
- Assume Mason Brewing brews 36,000 litres and 94,500 litres in 2018 and 2019, respectively.
C) Double Declining Balance method

3) What is the book value of the equipment at December 31st, 2019, under the three methods?
4) Mason Brewing decides to scrap the new micro-brew after reviewing 2019 results. On January 1, 2020 they
sell the Copper Kettle to a local buyer. Assuming Mason Brewing used the straight-line method, record
the disposal assuming a selling price of:
A) $ 30,000
B) $ 35,000
Answer #1
1) Recording the costs on June 30, 2018 to acquire the Copper Kettle.
Date Accounts Title Debit $ Credit $
30-Jun-18 Copper Kettle-Purchase 35340
Copper kettle - Freight in 830
Copper Kettle - Insurance 2800
Copper Kettle - Installation 630
Copper Kettle - Testing 1200
Cash 40800
(being purchase price, freight charges &
other expense on Copper kettle booked)
2) Depreciation of Copper Kettle:
A) Straight line:
For 2018 (40800-3000)*0.50/10 = 1890
For 2019 (40800-3000)/10 = 3780
B) Units of Production:
For 2018 (40800-3000)*36000/900000 = 1512
For 2019 (40800-3000)*94500/900000 = 3969
C) Double Declining Balance:
For 2018 40800 * (10% *2) / 2 = 4080
For 2019 (40800-4080) * (10% *2) = 7344
3) Book value of the equipment ie. Copper Kettle:
On Dec 31, 2019:
Straight line 40800 - 5670 = 35130
Units of production 40800 - 5481 = 35319
Double Declining 40800 - 11424 = 29376
4) Journal entries on disposal :
A) Selling Price = $30000
Date Accounts Title Debit $ Credit $
1-Jan-20 Cash 30000
Acc. Dep. - Copper Kettle 5670
Loss on disposal 5130
Copper Kettle 40800
(being loss on disposal of equipment booked)
B) Selling Price = $35000
1-Jan-20 Cash 35000
Acc. Dep. - Copper Kettle 5670
Loss on disposal 130
Copper Kettle 40800
(being loss on disposal of equipment booked)